# Retrieve to Explain: Evidence-driven Predictions with Language Models

### v2

@@ 3.3 R2E Retriever
The MLM in Section 3.2 has text embeddings that are trained to have high vector similarity with the paired answer embeddings in the answer set. We reasoned that two text embeddings would therefore have high vector similarity if they permit a similar distribution over answers, i.e. if they were semantically similar with respect to this task. This makes the MLM particularly well-suited for identifying corpus passages that are relevant to the user query. We therefore used this MLM for the Retriever module of R2E.
We first used the MLM to embed each of the masked evidence passages in the evidence corpus $\mathcal{D}^{e}$, where $\mathcal{D}^{e}\subset\mathcal{D}$. We partitioned the evidence embeddings according to the masked answer entity they contained, and created distinct FAISS search indices (Johnson et al., 2019) for each.
At inference time, the user’s cloze-style query is encoded with the MLM, yielding $q$. The Retriever selects the $k$ evidence passages $[d^{e}_{1i},...,d^{e}_{ki}]\subset\mathcal{D}^{e}$ with the highest cosine similarity to $q$ from each answer $a_{i}$’s search index. For our experiments, $k=64$. The query and retrieved evidence for each answer are then the inputs to the Reasoner.
@@ 3.4 R2E Reasoner
3.4.1 Training Objective
We train the Reasoner with a binary cross entropy loss to differentiate positive examples ($L=1$) from negative examples ($L=0$), i.e. to learn $p(L=1|a_{i},d^{q})$ when taking an entity $a_{i}$ and masked query $d^{q}$ as input. Positive examples were constructed from pairs ($a_{p}$, $d^{q}$) in $\mathcal{D}^{q}\subset\mathcal{D}$, the data split of Reasoner training queries. For each positive example, a corresponding negative example ($a_{n}$, $d^{q}$) was constructed by uniformly sampling $a_{n}\in\mathcal{A}$. For each ($a_{i}$, $d^{q}$) pair, positive or negative, the Reasoner receives the top $k$ evidence passages $[d^{e}_{1i},...,d^{e}_{ki}]$ fetched by the Retriever from the retrieval corpus of $a_{i}$. Under this negative sampling scheme, the objective $p(L=1|a_{i},d^{q})$ is closely related to the MLM multiclass objective $p(a_{i}|d^{q})$ at optimality (Appendix F); however, unlike multinomial regression, sampling negative examples avoids retrieving evidence for every possible alternative answer during training.
3.4.2 Inference
At inference time, we generate $p(L=1|a_{i},q)\,\forall\,a_{i}\in\mathcal{A}$ to score and rank the full answer set for the cloze-style query $q$, using the evidence fetched for $q$. This requires $|\mathcal{A}|$ nearest neighbour searches and forward passes through the Reasoner; however, since retrieval and reasoning for each answer is independent of all other answers, the process can be fully parallelized subject to computational resources.
3.4.3 Architecture
The R2E Reasoner architecture is shown in Figure 2.
@@ 3.4.3 Architecture
Next, the evidence combiner $g:(\mathbb{R}^{h})^{k}\rightarrow[0,1]$ merges the resulting $k$ query-evidence embeddings and generates $p(L=1|a_{i},q)$. As there is no inherent ordering among the $k$ vectors, we use a set transformer (Lee et al., 2019).
Since the Reasoner sees the answer $a_{i}$ only indirectly via the evidence embeddings, we can also view the score as the probability that the query and evidence embeddings are discussing the same entity.
@@ 3.5 R2E Explanations
R2E provides explanations in the form of Shapley values (Shapley et al., 1953; Lundberg & Lee, 2017) attributing the model prediction back to pieces of retrieved evidence, such that they sum up to the overall model score for the query.
Multiple methods exist for rapidly approximating Shapley values on deep learning features (Lundberg & Lee, 2017). Defining each of the $k$ inputs to the evidence combiner as a distinct feature gives a relatively small feature space. As a result, it is tractable to use a simple permutation sampling approach to approximate Shapley values (see Appendix E).
During training, we replaced query-evidence features at random with a learned NULL embedding. In addition to acting as a regularizer (akin to dropout), introducing the NULL embedding during training ensured that the model could handle missing features robustly when estimating Shapley values. For each training example, a dropout rate was sampled in $\text{Uniform}(0,1)$.
@@ 3.6 Post-hoc Frequency Bias Correction
Many potential answer sets suffer from class imbalance. In drug target identification, some targets are significantly more well-studied than others. As a result, the learned model $p(a_{i}|q)$ can be strongly correlated with the prior $p(a_{i})$.
While bias can be informative (e.g. reflecting the fact that some targets are involved in more diseases than others) it can also be misleading (e.g. reflecting publishing trends rather than underlying biology). To flexibly control for bias, we therefore use a corrected output based on the frequency of each answer in the training corpus and parameterized by $c\in[0,1]$ (Appendix G). When $c=0$ the scores and rankings are unaltered; when $c=1$, the rankings reflect the pointwise mutual information (PMI) of the query and answer, inspired by the use of PMI in NLP co-occurrence statistics (Church & Hanks, 1990). In the results we report both uncorrected ($c=0$; R2E-uncor) and partially corrected ($c=0.5$; R2E-cor) rankings, with the latter value selected using a validation set (Appendix D). In Shapley value explanations, the bias correction can be represented as an additive feature.
@@ 4 Experiments and Results
We evaluate R2E performance on three datasets aligned with drug target identification, each involving prediction over Genes:
  • [itemsep=0.0pt,topsep=0pt,partopsep=0pt]
  • Held-out Biomedical Literature: Predicting masked genes in biomedical literature sentences taken from abstracts published after the publication period of the training data and retrieval corpus.
  • Gene Description Facts: Predicting masked genes in sentences containing human-curated information about the gene, based on gene descriptions provided by UniProt (Consortium, 2022).
  • Clinical Trial Outcomes: Retrospectively predicting success or failure in historical clinical trial data based on evidence published before the trials, using the disease indication and drug target (gene).
For Gene Description Facts and Clinical Trial Outcomes, we also construct Evidence Annotations datasets to evaluate the alignment of R2E explanations with expert reasoning. Specifically, we look at the strength of relationship between R2E Shapley values and GPT-4 (Achiam et al., 2023) binary annotations of whether each piece of explanatory evidence is relevant or irrelevant to the query. We validate GPT-4 annotations against human drug discovery expert annotations.
Evaluation dataset statistics are summarised in Appendix H, and information on accessing datasets in Appendix A.
Table 1: Held-out Biomedical Literature: Ranking metrics on a dataset consisting of one sentence per gene in Genes for each year of 2020, 2021, and 2022. MLM and R2E trained on data published prior to 2020. MCS, R2E-Uncorrected and R2E-Corrected also retrieved data published prior to 2020. R2E-Corrected-Updated retrieved up to the year before the publication year of the query sentence.
Metric Query year Baselines R2E
Freq MCS MLM Uncor Cor Cor-Updated
2020 $<$0.001 0.182 0.181 0.198 0.233 -
MRR 2021 $<$0.001 0.172 0.169 0.187 0.215 0.223
2022 $<$0.001 0.167 0.164 0.178 0.205 0.219
2020 7661 3280 3465 2803 2489 -
MR 2021 7834 3568 3789 3032 2695 2544
2022 7931 3770 4016 3287 2902 2623
2020 $<$0.001 0.268 0.269 0.291 0.333 -
h@10 2021 $<$0.001 0.251 0.252 0.274 0.313 0.324
2022 $<$0.001 0.243 0.243 0.260 0.295 0.312
2020 0.014 0.443 0.438 0.484 0.521 -
h@200 2021 0.014 0.422 0.416 0.456 0.497 0.509
2022 0.013 0.404 0.398 0.435 0.473 0.496
@@ 4.1 Metrics
For ranking Genes on Held-out Biomedical Literature and Gene Description Facts, we report mean reciprocal rank (MRR), mean rank (MR), hits@10 (h@10) and hits@200 (h@200). In the case of Gene Description Facts, we used macro ranking metrics to ensure each gene is given equal weight irrespective of frequency. For Clinical Trial Outcomes we report AUROC. For Evidence Annotations, we report AUROC for the R2E Shapley scores of evidence sentences against GPT-4 annotations, and accuracy when validating GPT-4 against human expert annotations.
4.2 Baselines
@@
In addition to MLM, we include two baselines throughout: FREQ and MCS. For FREQ, entities were scored according to their frequency in the training set of $\mathcal{D}$. For MCS (mean cosine similarity), each entity $a_{i}$ was scored by computing $\frac{1}{64}\sum_{j=1}^{64}(d^{e}_{ji}\cdot q)/(\|d^{e}_{ji}\|\|q\|)$ for the query $q$.
For Clinical Trial Outcomes, we also include a genetics baseline, described in detail in Appendix P.
4.3 Predicting Genes in Held-out Biomedical Literature
4.3.1 Dataset Construction
For all experiments in this section, we trained the MLM (R2E Retriever) and R2E Reasoner only on biomedical literature data published prior to 2020. Except where specified, R2E also only retrieved data published prior to 2020 (Figure S3). We then constructed a Held-out Biomedical Literature evaluation dataset from publicly-available paper abstracts. We generated a balanced dataset by sampling one sentence per unique gene in Genes for each of the years 2020, 2021, and 2022. See Appendix B for more detail.
4.3.2 Results
R2E improved on the baselines over all year subsets, both with and without bias correction (Table 1). Bias-corrected R2E improved on uncorrected performance, consistent with the use of a balanced evaluation dataset. We show results on an imbalanced dataset (without stratification by gene in Genes) in Appendix I, demonstrating comparable performance between R2E, MCS and MLM.
To test R2E’s ability to leverage retrieved literature that it was not trained on, we enabled retrieval up to the year preceding the query sentence publication (rather than strictly prior to 2020). This improved performance (Table 1).
We performed two ablation studies: replacing the R2E Retriever with a pre-trained PubMedBERT model (Gu et al., 2021), and replacing the query-evidence encoder of the R2E Reasoner with a Hadamard product $q\circ d^{e}$. R2E outperformed both ablations (Appendix J).
We additionally investigated restricting retrieval to literature data published prior to the query sentence during training, with little consistent impact on performance (Appendix K).
4.4 Predicting Genes in Gene Description Facts
4.4.1 Dataset Construction
@@
We extracted descriptions of protein functions for our Genes entity set from UniProt (Universal Protein Resource) (Consortium, 2022). Each description is a human-written summary of function, and consists of one or more sentences. We used a combination of regular expressions and prompting the Claude 2.0 language model from Anthropic [^1] https://www.anthropic.com/ to extract a set of [MASK]-containing facts from each description. Further details of the source and preprocessing of the dataset, including the Claude prompt and an example description and facts, are found in Appendix L. R2E was trained on, and retrieved from, all years of literature evidence for the Gene Description Facts evaluation.
@@
We also constructed an Evidence Annotations dataset by having GPT-4 annotate (as query relevant or irrelevant) all evidence for 50 randomly sampled Gene Description Facts query-entity pairs (positive examples), and the same 50 queries with randomly sampled alternative entities (negative examples). See Appendix M for the prompting strategy and Appendix N for illustrative annotations. A human drug discovery expert annotated all evidence sentences for a subset of 8 randomly sampled examples (4 positive, 4 negative), following equivalent instructions to the GPT-4 prompt, to validate GPT-4 annotations.
4.4.2 Results
R2E substantially improved on all baselines, both with and without bias correction (Table 2). Bias correction again outperformed the uncorrected model. R2E metrics appear to show greater improvement over baselines than for the Held-out Biomedical Literature dataset. Additionally, there was a strong correlation between evidence Shapley values and GPT-4 relevance annotations, with an AUROC of 0.824. GPT-4 had an accuracy of 0.715 when evaluated against the human-expert annotations.
Table 2: Gene Description Facts: R2E macro ranking metrics on cloze-style queries adapted from human-curated gene descriptions.
@@ 4.4.2 Results
4.5 Predicting Clinical Trial Outcomes
4.5.1 Dataset Construction
We constructed a benchmark of gene-disease pairs (therapeutic hypotheses) with success or failure outcomes in clinical trials as per Nelson et al. 2015 (Appendix O), using the PharmaProjects database (Citeline, ). This benchmark focuses on in vivo efficacy of therapeutic hypotheses as demonstrated by the transition of drugs from Phase II/III clinical trials to regulatory approval.
As genetics data is widely acknowledged to predict successful clinical development (Nelson et al., 2015; Ochoa et al., 2022), we use the most recently published PharmaProjects-aligned genetics dataset (Minikel et al., 2023) (Appendix P) as a competitive baseline. In order to validate our Clinical Trial Outcomes data, we corroborated the published result that the probability of clinical success of therapeutic hypotheses supported by genetics evidence is approximately double the probability without supporting genetics evidence (Relative Success: 1.98; 95% CI (1.76, 2.24)).
We also constructed an Evidence Annotations dataset by having GPT-4 assess the relevance of all evidence for 100 Clinical Trial Outcome therapeutic hypotheses associated with success (50 success, 50 failure; randomly sampled). See Appendix Q for the prompting strategy and Appendix R for illustrative annotations. A human drug discovery expert annotated all evidence sentences for the 8 therapeutic hypotheses for which they had most specialised knowledge (4 success, 4 failure), following equivalent instructions to the GPT-4 prompt, to validate GPT-4 annotations.
To avoid temporal leakage due to the reporting of clinical trial results in the literature, we trained the Reasoner and Retriever with pre-2005 literature (Section 3.1) and filtered out drugs with dated evidence prior to 2005 (Appendix O). We predicted therapeutic hypotheses using the query template “[MASK] is a promising drug target for the treatment of {DISEASE}.”, substituting {DISEASE} with the disease annotated in PharmaProjects (Appendix S).
4.5.2 Multimodality via Templating into Natural Language
We assessed R2E’s ability to reason from genetics data by generating a sentence for every row in the genetics dataset (77,645 total), with the simple template “[MASK] is genetically associated with {MeSH name}.”. The MeSH name, as supplied in Minikel et al. 2023, was programmatically reformatted to better align with naming conventions in the biomedical literature (details in Appendix S). The genetics corpus was given to the R2E Retriever alone and in combination with the year-split biomedical literature.
4.5.3 Results
Table 3: Clinical Trial Outcomes: AUROC for R2E retrieving from literature-alone, genetics-alone, or both; in comparison to baselines.
Model Modality AUROC

### v4

@@ 3.3 R2E Retriever
The MLM in Section 3.2 produces text embeddings that are trained to have a high inner-product with the paired answer embeddings in the answer set. We reasoned that two text embeddings would therefore have high similarity if they permit a similar distribution over answers, i.e. if they were semantically similar with respect to this task. This makes the MLM well-suited to identifying corpus passages that are relevant to the user query and so we used this MLM as the R2E Retriever.
We first used the MLM to embed each of the masked evidence passages in the evidence corpus $\mathcal{D}^{e}$, where $\mathcal{D}^{e}=\mathcal{D}^{\mathrm{MLM}}$ for Reasoner training (Section 3.4; typically $D^{e}=\mathcal{D}$ at inference). We partitioned evidence embeddings by the masked answer entity they contained, and created distinct FAISS search indices Johnson et al. (2019) for each to enable efficient answer-specific retrieval.
At inference time, the user’s cloze-style query $q$ is encoded with the MLM. The Retriever selects $k$ evidence passages $[d^{e}_{1i},...,d^{e}_{ki}]\subset\mathcal{D}^{e}$ with the highest cosine similarity to $q$ from each answer $a_{i}$’s search index (we use $k=64$). The query embedding and retrieved evidence embeddings for each answer are the inputs to the Reasoner.
@@ 3.4 R2E Reasoner
Training Objective
We train the Reasoner with a binary cross entropy loss to differentiate positive examples ($L=1$) from negative examples ($L=0$), i.e. to learn $p(L=1|a_{i},d^{q})$ when taking an entity $a_{i}$ and masked query $d^{q}$ as input, where $d^{q}\in\mathcal{D}^{q}\subset\mathcal{D}\setminus\mathcal{D}^{e}$ ($\mathcal{D}^{e}$ excluded to avoid trivial inference by retrieving $d^{q}$ from $\mathcal{D}^{e}$). Positive examples were constructed from pairs $(a_{p},d^{q})\forall d^{q}\in\mathcal{D}^{q}$, where $a_{p}$ is the true masked answer in $d^{q}$. For each positive example, a corresponding negative example $(a_{n},d^{q})$ was constructed by uniformly sampling $a_{n}\in\mathcal{A}\backslash\{a_{p}\}$. For each $(a_{i},d^{q})$ pair, positive or negative, the Reasoner receives the top $k$ evidence passages $[d^{e}_{1i},...,d^{e}_{ki}]$ fetched by the Retriever from the retrieval corpus of $a_{i}$. For negatives, due to sampling of $a_{n}$, retrieved evidence corresponds to a different entity to the answer entity masked in the query. Under this sampling scheme, the objective $p(L=1|a_{i},d^{q})$ is closely related to the MLM multiclass objective $p(a_{i}|d^{q})$ at optimality (Appendix F); however, unlike multinomial regression, sampling avoids needing to retrieve evidence for all answers for each training example.
Inference
At inference time, we use $p(L=1|a_{i},q)\,\forall\,a_{i}\in\mathcal{A}$ to score and rank the full answer set for the cloze-style query $q$, using the evidence fetched for $q$. This requires $|\mathcal{A}|$ nearest neighbour searches and forward passes through the Reasoner; however, since retrieval and reasoning independent for each answer, the process can be fully parallelized subject to computational resources. See Appendix H for profiling of inference speeds.
Architecture
@@
Next, the evidence combiner $g:(\mathbb{R}^{h})^{k}\rightarrow[0,1]$ generates $p(L=1|a_{i},q)$ from the $k$ query-evidence embeddings. There is no inherent ordering among the $k$ vectors, so we use a set transformer Lee et al. (2019).
Answers are masked in both the query and answer-specific evidence so that the Reasoner sees $a_{i}$ only indirectly via evidence embeddings. As a result, the score reflects the probability that query and evidence embeddings relate to the same entity.
@@ 3.5 R2E Explanations
R2E provides explanations in the form of Shapley values Shapley et al. (1953); Lundberg and Lee (2017) - the average expected marginal contribution of each piece of evidence to the overall model score for the query. Shapley values enable attribution of the model prediction back to pieces of retrieved evidence, such that they sum up to the overall score.
Multiple methods exist for rapidly approximating Shapley values on deep learning features Lundberg and Lee (2017). Defining each of the $k$ inputs to the evidence combiner as a distinct feature gives a relatively small feature space, making it tractable to use a permutation sampling approach to approximate Shapley values. See Appendix E for the full algorithm and Appendix H for profiling.
During training, we replaced query-evidence features at random with a learned NULL embedding. In addition to acting as a regularizer (akin to dropout), introducing the NULL embedding during training ensured that the model could handle missing features robustly when estimating Shapley values. For each training example, the evidence dropout rate was sampled in $\text{Uniform}(0,1)$.
@@ 3.6 Post-hoc Frequency Bias Correction
Many answer sets suffer from class imbalance. In drug target identification, some targets are significantly more well-studied than others. As a result, the learned model $p(a_{i}|q)$ can be strongly correlated with the prior $p(a_{i})$.
While such bias can be informative (e.g. reflecting the fact that some targets are involved in more diseases than others) it can also be misleading (e.g. reflecting publishing trends rather than underlying biology). To flexibly control for bias, we introduce a method to correct the model output score based on the frequency of answers in the training corpus, resulting in an up-ranking of less frequently mentioned answers, as detailed in Appendix G. The correction is parameterized by $c\in[0,1]$: when $c=0$ the scores and rankings are unaltered; when $c=1$, the rankings reflect the pointwise mutual information (PMI) of the query and answer, inspired by PMI use in NLP co-occurrence statistics Church and Hanks (1990). In the results we report both uncorrected ($c=0$; R2E-uncor) and partially corrected ($c=0.5$; R2E-cor; selected using validation set, Appendix D) rankings. In Shapley value explanations, the bias correction can be represented as an additive feature.
@@ 4 Experiments and Results
We evaluate R2E performance on three datasets aligned with drug target identification, which we publicly release (Appendix A):
  • Held-out Biomedical Literature: Predicting masked genes in biomedical literature sentences from abstracts published after the publication of the training data and retrieval corpus.
  • Gene Description Facts: Predicting masked genes in sentences from human-curated gene descriptions adapted from UniProt Consortium (2022).
  • Clinical Trial Outcomes: Retrospectively predicting success or failure in clinical trials based on evidence published before the trials, using the disease indication and drug target (gene).
For Gene Description Facts and Clinical Trial Outcomes, we also construct Evidence Annotations datasets to evaluate the alignment of R2E explanations with expert reasoning. We look at the strength of relationship between R2E Shapley values and GPT-4 Achiam et al. (2023) binary annotations of whether each piece of explanatory evidence is relevant or irrelevant to the query. We validate GPT-4 annotations against human expert annotations. More detailed usability testing of R2E Shapley values is left to future work.
For dataset summary statistics see Appendix I.
@@ 4.1 Metrics
For ranking Genes on Held-out Biomedical Literature and Gene Description Facts, we report mean reciprocal rank (MRR), mean rank (MR), hits@10 (h@10) and hits@200 (h@200). For Gene Description Facts, we used macro metrics to give each gene equal weight irrespective of frequency. For Clinical Trial Outcomes we report AUROC, and include relative success results in Appendix U for consistency with Minikel et al. (2024). We compare AUROCs using DeLong test, and relative successes using Z-test, reporting confidence intervals using Katz method Katz et al. (1978). For Evidence Annotations, we report AUROC for the R2E Shapley scores of evidence sentences against GPT-4 annotations, and accuracy when validating GPT-4 against human expert annotations.
4.2 Baselines and Ablations
@@
In addition to MLM (Section 3.2), we include two baselines throughout: FREQ and MCS. For FREQ, entities were scored according to their frequency in the training set of $\mathcal{D}$. For MCS (mean cosine similarity), each entity $a_{i}$ was scored by computing $\frac{1}{64}\sum_{j=1}^{64}(d^{e}_{ji}\cdot q)/(\|d^{e}_{ji}\|\|q\|)$ for the query $q$.
For Clinical Trial Outcomes, we include a competitive genetics baseline used throughout the pharmaceutical industry (in-depth setup in Appendix Q). Alongside other relative success results in Appendix U, we compare to a few-shot, chain-of-thought, retrieval-augmented GPT-4 baseline (setup in Appendix V). For extensive ablations of R2E, including the Retriever, Reasoner and literature bias correction, see Appendix L.
4.3 Held-out biomedical literature
Given their greater orthogonality to the R2E training objective, we choose to focus on Gene Description Facts and Clinical Trial Outcomes in the main text, and save complete results for Held-out Biomedical Literature for Appendix J (Table 5). For the latter, R2E outperformed all baselines and was able to leverage retrieved literature that it was not trained on, further improving performance.
4.4 Gene Description Facts
Dataset Construction
@@ 4.4 Gene Description Facts
We sought to validate that R2E could perform well on predicting protein-coding genes in human-curated descriptions of gene function. We extracted descriptions of protein functions for our Genes entities from UniProt (Universal Protein Resource) Consortium (2022). Each UniProt description is a human-written summary of a protein’s function, and consists of one or more sentences. We used a combination of regular expressions and Anthropic’s Claude 2.0 to extract [MASK]-containing facts from each description. Further details of the source and preprocessing of the dataset, including the Claude prompt and an example gene description with extracted facts, are found in Appendix M. R2E was trained on, and retrieved from, all years of literature evidence for the Gene Description Facts evaluation.
@@
We also constructed an Evidence Annotations dataset by having GPT-4 (prompt in Appendix N) annotate as query-relevant or irrelevant, all evidence for 50 randomly sampled Gene Description Facts query-entity pairs (positive examples), and the same 50 queries with randomly sampled alternative entities (negative examples), obtaining 6400 annotated query-evidence pairs. To validate GPT-4 annotations, a human drug discovery expert following the GPT-4 prompt annotated all 512 query-evidence pairs for a subset of 8 randomly sampled examples (4 positive, 4 negative).
Results
R2E substantially improved on all baselines, both with and without bias correction (Table 1). As expected, bias correction was helpful. R2E metrics here appear to show greater improvement over baselines than for the Held-out Biomedical Literature dataset in Table 5. This may reflect a tendency for gene descriptions to describe more well-established knowledge than literature; as a result, R2E may benefit from its access to such facts, when more directly stated in the retrieved evidence sentences.
Additionally, there was a strong correlation between evidence Shapley values and GPT-4 relevance annotations (AUROC: 0.824). See Appendix O for a case study of examples. Combined with a $71.5$% agreement rate between GPT-4 and human-expert annotations, the agreement between R2E and GPT-4 suggests that R2E has correctly learnt to prioritise evidence for its predictions.
Table 1: Gene Description Facts: R2E macro ranking metrics.
@@ Results
4.5 Clinical Trial Outcomes
Dataset Construction
We constructed a benchmark of gene-disease pairs (therapeutic hypotheses) from clinical trials as per Nelson et al. 2015, using the PharmaProjects database Citeline (1,449 success, 4,222 failure, Appendix P). This benchmark focused on in vivo efficacy of therapeutic hypotheses as demonstrated by the transition of drugs associated with such hypotheses from Phase II/III clinical trials to regulatory approval.
To avoid leakage due to reporting of clinical trials in the literature, we removed drugs investigated prior to 2005 (Appendix P) and used pre-2005 literature for R2E training and retrieval (Appendix B). We scored therapeutic hypotheses using a query template “[MASK] is a promising drug target for the treatment of {DISEASE}.", substituting in the PharmaProjects disease (Appendix T).
As the ability of genetics methods such as locus-to-gene Mountjoy et al. (2021) to predict successful clinical development Nelson et al. (2015); Ochoa et al. (2022); Minikel et al. (2024) drives their wide use in target identification, we used the most recently published PharmaProjects-aligned dataset of genetics predictions Minikel et al. (2024) (Appendix Q) as a competitive baseline. In order to validate our Clinical Trial Outcomes data, we corroborated the published result Minikel et al. (2024) that the probability of clinical success of therapeutic hypotheses supported by genetics evidence is approximately double the probability without supporting genetics evidence (relative success: 1.98; 95% CI (1.76, 2.24); Appendix U.2).
We also constructed an Evidence Annotations dataset with GPT-4 (prompt in Appendix R) assessing the relevance of all 64 evidence passages for 100 Clinical Trial Outcome therapeutic hypotheses (50 success, 50 failure; randomly sampled), obtaining 6400 annotated query-evidence pairs. To validate GPT-4 annotations, a human drug discovery expert following the GPT-4 prompt annotated all 512 query-evidence pairs for a subset of 8 hypotheses for which they had most knowledge (4 success, 4 failure).
Multimodality via Templating into Natural Language
We assessed R2E’s ability to reason from genetics by generating a sentence for every row in the genetics dataset used in the genetics baseline (77,645 total), with the simple template “[MASK] is genetically associated with {MeSH name}.”. The MeSH name, as supplied in Minikel et al. 2024, was programmatically reformatted to better align with naming conventions in the biomedical literature (details in Appendix T). This genetics corpus was given to the R2E Retriever alone and in combination with the pre-2005 biomedical literature.
Table 2: Clinical Trial Outcomes: AUROC for R2E with retrieval corpus of literature-alone, genetics-alone, or both combined. For relative success metrics, including comparison to a few-shot, chain-of-thought, retrieval-augmented GPT-4 baseline, see Figure 4 and Appendix U.
Model Corpus AUROC